# Conservative Closures of the Vlasov-Poisson Equations Discretized with a Symmetrically Weighted Hermite Spectral Expansion in Velocity

### v1

@@
Conservative Closures of the Vlasov-Poisson Equations Discretized with a Symmetrically Weighted Hermite Spectral Expansion in Velocity
Opal Issan oissan@ucsd.edu Oleksandr Koshkarov Federico D. Halpern Boris Kramer Gian Luca Delzanno Department of Mechanical and Aerospace Engineering, University of California San Diego, La Jolla, CA, USA T-5 Applied Mathematics and Plasma Physics Group, Los Alamos National Laboratory, Los Alamos, NM, USA General Atomics, P.O. Box 85608, San Diego, CA, USA
@@ 1 Introduction
Most of the spectral methods for the kinetic equations have focused on Hermite functions in velocity space weighted by a Gaussian exponential function [23, 17, 31, 25, 9]. This idea originated from Grad’s 1949 paper [13, §4], which proposed a 13-moment equation coupled with the closure by truncation (assuming the 14th moment is zero). The Hermite spectral expansion is particularly advantageous since it can approximate near-Maxwellian distributions with relatively few basis functions. However, the proposed 13-moment equations lack global hyperbolicity due to complex eigenvalues in the flux Jacobian operator. It is important to preserve the hyperbolic nature of the equations to avoid introducing unphysical dissipation into the strictly advective collisionless system. Since then, several works have introduced a regularized moment system to preserve hyperbolicity [6, 33, 22].
In this short note, we investigate alternative conservative closures for the one-dimensional collisionless Vlasov-Poisson equations discretized in velocity via the symmetrically weighted (SW) Hermite spectral expansion [19, 17]. The primary benefit of the SW Hermite spectral expansion is its numerical stability properties, and its major drawback is the lack of conservation properties and fluid-kinetic coupling since all the Hermite spectral coefficients are needed to describe the fluid quantities. We focus on closures that conserve mass, momentum, or energy. Additionally, we examine the proposed conservative closures preservation of hyperbolicity, and anti-symmetry (important for stability purposes) of the Vlasov equation. We verify numerically the analytically derived conservative closures on a classic benchmark problem: the Langmuir wave.
This short note is organized as follows. First, section 2 briefly reviews the one-dimensional collisionless Vlasov-Poisson equations and their velocity discretization via the SW Hermite spectral expansion. Section 3 presents three conservative closures: (1) closure by truncation, which we previously rigorously analyzed in [19], (2) closure by momentum, and (3) closure by energy. Section 4 discusses the closures’ ability to preserve the Vlasov equation hyperbolicity and anti-symmetric structure. We test numerically the analytic derivations of the closure properties on the Langmuir wave test in section 5. Lastly, in section 6, we discuss conclusions and further considerations.
@@ 2 Vlasov-Poisson Equations: Symmetrically Weighted Hermite Spectral Expansion in Velocity
where $E(x,t)$ is the normalized electric field, $m^{s}$ is the normalized mass of species $s$, and $q^{s}$ is the normalized charge of species $s$. All quantities in the Vlasov-Poisson equations (1)–(2) are normalized in the centimeter-gram-seconds system of units as follows:
@@
The velocity coordinate $v$ is shifted by $u^{s}\in\mathbb{R}$ and scaled by $\alpha^{s}\in\mathbb{R}_{>0}$ [34], which are two tunable parameters specified by the user. The SW Hermite basis function of degree $n\in\mathbb{N}_{0}$ is
@@
$\frac{\partial C_{n}^{s}(x,t)}{\partial t}$ $+\frac{\partial}{\partial x}\left(\alpha^{s}\sqrt{\frac{n}{{2}}}C_{n-1}^{s}(x,t)+\alpha^{s}\sqrt{\frac{n+1}{2}}C_{n+1}^{s}(x,t)+u^{s}C_{n}^{s}(x,t)\right)$ (5)
$+\frac{q^{s}}{m^{s}\alpha^{s}}E(x,t)\left(-\sqrt{\frac{n}{2}}C_{n-1}^{s}(x,t)+\sqrt{\frac{n+1}{2}}C_{n+1}^{s}(x,t)\right)=0,$
@@
Note that the SW Hermite closure introduces an error in both the advection $(v\partial f^{s}/\partial x)$ and acceleration $(E\partial f^{s}/\partial v)$ terms. The Poisson equation (2) after Galerkin projection is given by
@@ 3 Conservative Closures for the Symmetrically Weighted Hermite Formulation
The main question we seek to answer is: Is there a closure in the form of Eq. (6) that improves the mass, momentum, energy, or Casimir (e.g., $\mathcal{L}_{2}$ norm) conservation properties of the SW Hermite formulation? We begin by deriving the mass, momentum, and energy fluid equations and their respective drift rates with an arbitrary closure. Then, we set the drift rates to zero and derive a conservative closure. The proposed closures are (1) closure by truncation, (2) closure by momentum, and (3) closure by energy. The closures result in a system of partial differential-algebraic equations for the coefficients motivated by the fluid equations, which are the first three moments with respect to the velocity monomials $,1,v,v^{2},$ of the Vlasov equation (1).
@@ 3.1 Mass Conservative Closure (Equivalent to Closure by Truncation)
Recall that we use periodic boundary conditions. Therefore, the closure cannot impact the mass drift rate if $N_{v}$ is even. However, if $N_{v}$ is odd, then setting $\int_{0}^{\ell}E(x,t)C_{N_{v}}^{s}(x,t)\mathrm{d}x=0$ causes the total mass drift rate to vanish. The closure by truncation, i.e. $C_{N_{v}}^{s}(x,t)=0$, is one example of a closure resulting in a total mass conservative formulation however, it is the only closure that satisfies the continuity equation.
@@ 3.2 Momentum Conservative Closure
Therefore, the fluid momentum equation is satisfied and the momentum drift rate vanishes if and only if
$C_{N_{v}}^{s}(x,t)=C_{N_{v}-1}^{s}(x,t)\begin{cases}\frac{\alpha^{s}\sqrt{2N_{v}}}{u^{s}}\qquad&\text{if }N_{v}\text{ is odd},\\ \frac{u^{s}}{\alpha^{s}\sqrt{2N_{v}}}\qquad&\text{if }N_{v}\text{ is even}.\\ \end{cases}$ (15)
We refer to this closure as closure by momentum. For both cases of $N_{v}$ being odd or even, the closure term $C_{N_{v}}^{s}(x,t)$ is proportional to the last Hermite expansion coefficient $C_{N_{v}-1}^{s}(x,t)$. It is also important to mention that this closure requires the velocity shifting parameter $u^{s}\neq 0$ for all species $s$, which can be easily imposed by performing a Galilean transformation. In the case of $N_{v}$ being even and $u^{s}=0$, the closure by truncation also satisfies the momentum equation and thus preserves the total momentum.
@@ 3.3 Energy Conservative Closure
where $\mu^{s}_{n}(x,t)=\frac{q^{s}}{m^{s}}\int_{0}^{x}E(y,t)C_{n}^{s}(y,t)\mathrm{d}y$. The extra term involving $\mu_{n}^{s}(x,t)$ stems from integrating the SW Poisson equation (7) to arrive at the SW Ampere’s equation. The total energy drift rate in Eq. (17) is zero,$\frac{\mathrm{d}\mathcal{E}}{\mathrm{d}t}=0$, if and only if
@@ 4 Closure Structure Preservation: Hyperbolicity and Anti-symmetry
where $\sigma^{s}\in\mathbb{R}$ is an arbitrary scalar component of the closure function relating linearly $C_{N_{v}}^{s}(x,t)$ to $C_{N_{v}-1}^{s}(x,t)$. Similarly, by Eq. (1), the source operator with a general closure in the form of Eq. (6) is given by
@@ 4.1 Hyperbolicity
So far we have treated the case with constant $\alpha^{s}$ and $u^{s}$. However, there is a strong mathematical and physical motivation to adapt spatially and temporally the Hermite shifting and scaling parameters, denoted as $u^{s}$ and $\alpha^{s}$ in Eq. (4). The mathematical considerations stem from improving the convergence of the Hermite spectral discretization. The physical motivation stems from the Hermite basis function in Eq. (4) resembling a Maxwellian distribution (thermodynamic equilibrium), such that $u^{s}$ mimics the characteristic mean flow and $\alpha^{s}$ mimics the thermal velocity. Therefore, the Hermite parameters can be adapted in space and time based on the evolution of such macroscopic quantities. An adaptive approach based on evolving the macroscopic quantities was first proposed in the seminal paper by Grad [13], which suffers from a lack of hyperbolicity. On the other hand, recent work by [29] showed that the adaptivity in time of the asymmetrically weighted (AW) Hermite parameters (i.e. $u^{s}$ and $\alpha^{s}$) does not alter the governing equations of the expansion coefficients and relies on projecting the expansion coefficients at a fixed instant in time of the Hermite parameters. This allows the adaptive in-time AW and SW formulations to remain hyperbolic. The adaptivity in space of the Hermite parameters introduces new terms in the form of $C_{n}^{s}(x,t)\frac{\partial}{\partial x}\alpha^{s}(x)$ and $C_{n}^{s}(x,t)\frac{\partial}{\partial x}u^{s}(x)$ as additional “source” terms. These terms do not directly affect the Jacobian flux matrix and similar to time-adaptivity require a Hermite projection step onto the new basis at a fixed instant in time [24]. Therefore, by following the proposed adaptive approach for the AW Hermite method in [29, 24], a space and time adaptive approach of the SW Hermite method also preserves the system’s hyperbolicity.
@@ 4.2 Anti-symmetry
Anti-symmetry (also known as skew-symmetry) is an additional important structure for a spectral method to preserve. Aside from capturing the correct dynamics, an anti-symmetric discretization guarantees unconditional numerical stability [8, 19]. Our previous work [19, §3.2] explains in detail the anti-symmetric structure of the Vlasov advection operator. In short, it requires the flux Jacobian matrix in Eq. (20) to be symmetric, i.e. $\mathbf{J}^{s}=(\mathbf{J}^{s})^{\top}$, and the source operator in Eq. (21) to be anti-symmetric, i.e. $\mathbf{S}^{s}(x,t)=-\mathbf{S}^{s}(x,t)^{\top}$. As previously stated in [19, §3.2.2], anti-symmetry is only preserved via the closure by truncation, otherwise $\mathbf{S}^{s}(x,t)\neq-\mathbf{S}^{s}(x,t)^{\top}$. Additionally, the adaptivity in time of the Hermite parameters following the work by [29] does not change the governing equations and will preserve anti-symmetry for the closure by truncation. However, adaptivity in space breaks anti-symmetry by introducing non-zero terms on the main diagonal of the source operator.
@@ 5 Numerical Results
How does the choice of closure influence the conservation properties? Figure 1 shows the total mass, momentum, energy, and $\mathcal{L}_{2}$ norm conservation results with $N_{v}=\{10,11\}$ for the three proposed closures: (1) closure by truncation, (2) closure by momentum, and (3) closure by energy. The numerical results agree with the analytical derivations in section 3. Closure by truncation can conserve the $\mathcal{L}_{2}$ norm of the particle distribution function and total mass if $N_{v}$ is odd, closure by momentum can only conserve total momentum, and closure by energy can only conserve total energy. Additionally, Figure 1 indicates that the magnitude of the drift rate is comparable for the three closures, which starts to plateau at around $t\sim 25$. The closure by momentum and energy with $N_{v}=10$ become unstable before the final timestamp. Recall that numerical stability properties are guaranteed only for the closure by truncation since it preserves anti-symmetry.
How does the choice of closure impact the electric field? Figure 2 compares the electric field with $N_{v}=11$ Hermite discretization for the three closures. The results for the closure by momentum and closure by energy are polluted by high wavenumber oscillations in comparison to the closure by truncation. This suggests that the closure by truncation may be the most appropriate conservative closure for the SW Hermite method.
(a) closure by truncation w/ $N_{v}=10$ (even)
(b) closure by truncation w/ $N_{v}=11$ (odd)
(c) closure by momentum in Eq. (15) $N_{v}=10$ (even)
(d) closure by momentum in Eq. (15) w/ $N_{v}=11$ (odd)
(e) closure by energy in Eq. (18) w/ $N_{v}=10$ (even)
(f) closure by energy in Eq. (18) $N_{v}=11$ (odd)
Figure 1: Langmuir wave conservation results with the (a/b) closure by truncation, (c/d) closure by momentum, and (e/f) closure by energy. The numerical results verify the analytic derivations in Eqns. (11), (14), and (17).

### v2

# Conservative Closures of the Vlasov-Poisson Equations Based on Symmetrically Weighted Hermite Spectral Expansion

@@
Abstract
We derive conservative closures for the Vlasov-Poisson equations discretized in velocity via the symmetrically weighted Hermite spectral expansion. We demonstrate that no closure can simultaneously restore the conservation of mass, momentum, and energy in this formulation. The properties of the analytically derived conservative closures of each conserved quantity are validated numerically by simulating an electrostatic benchmark problem: the Langmuir wave. Both the numerical results and analytical analysis indicate that closure by truncation (i.e. setting the last Hermite moment to zero) is the most suitable conservative closure for the symmetrically weighted Hermite formulation.
@@ 1 Introduction
Most spectral methods for kinetic equations have focused on using Hermite functions in velocity space, weighted by a Gaussian exponential function [23, 17, 31, 25, 9, camporeale_sps_2016]. The Hermite spectral expansion is particularly effective because it can approximate near-Maxwellian distributions with relatively few basis functions. There are mainly two classes of Hermite discretization. The first, introduced by Grad in 1949 [13, §4], evolves the Hermite basis dynamically in space and time by its dependence on fluid velocity and temperature. In this method, conservation of mass, momentum, and energy is enforced directly by solving the fluid equations, independently of the closure. The closure only impacts hyperbolicity, which must be maintained to avoid unphysical dissipation in the strictly advective collisionless system [6, 33, 22]. The second approach uses a fixed Hermite basis through either an asymmetrically weighted (AW) or symmetrically weighted (SW) expansion. The AW expansion conserves mass, momentum, and energy regardless of the closure term due to its fluid-kinetic coupling, where the first few Hermite coefficients describe the fluid moments [31]. Its main drawback is that it suffers from numerical instability. Conversely, the SW expansion is numerically stable but lacks conservation properties and fluid-kinetic coupling, such that all Hermite coefficients are required to describe the fluid quantities [31, 17, 19]. Therefore, the closure of the SW expansion impacts its conservation properties.
In this work, we investigate alternative conservative closures for the one-dimensional collisionless Vlasov-Poisson equations discretized in velocity via the SW Hermite spectral expansion [17, 19]. We derive new closures that restore the SW conservation of total mass, momentum, or energy. Additionally, we examine the proposed conservative closures preservation of hyperbolicity, and anti-symmetry (important for stability purposes) of the Vlasov equation. We verify numerically the analytically derived conservative closures on a classic benchmark problem: the Langmuir wave.
This paper is organized as follows. First, section 2 briefly reviews the one-dimensional collisionless Vlasov-Poisson equations and their velocity discretization via the SW Hermite spectral expansion. Section 3 presents three conservative closures: (1) closure by truncation, which we previously rigorously analyzed in [19], (2) closure by momentum, and (3) closure by energy. Section 4 discusses the closures’ ability to preserve the Vlasov equation hyperbolicity and anti-symmetric structure. We test numerically the analytic derivations of the closure properties on the Langmuir wave test in section 5. Lastly, in section 6, we discuss conclusions and further considerations.
@@ 2 Vlasov-Poisson Equations: Symmetrically Weighted Hermite Spectral Expansion in Velocity
where $E(x,t)$ is the electric field, $m^{s}$ is the mass of species $s$, and $q^{s}$ is the charge of species $s$. All quantities in the Vlasov-Poisson equations (1)–(2) are normalized in the centimeter-gram-seconds system of units as follows:
@@
The velocity coordinate $v$ is shifted by $u^{s}\in\mathbb{R}$ and scaled by $\alpha^{s}\in\mathbb{R}_{>0}$ [34], which are two tunable parameters specified by the user. The Hermite parameters $u^{s}$ and $\alpha^{s}$ correspond to the characteristic mean flow and thermal velocity for each species $s$, as the zeroth-order SW Hermite basis functions $\psi_{0}(\xi_{s})$ resemble a Maxwellian distribution. These parameters are typically set based on the initial condition’s characteristic mean flow and thermal velocity. The SW Hermite basis function of degree $n\in\mathbb{N}_{0}$ is
@@
$\frac{\partial C_{n}^{s}(x,t)}{\partial t}$ $+\underbrace{\frac{\partial}{\partial x}\left(\alpha^{s}\sqrt{\frac{n}{{2}}}C_{n-1}^{s}(x,t)+\alpha^{s}\sqrt{\frac{n+1}{2}}C_{n+1}^{s}(x,t)+u^{s}C_{n}^{s}(x,t)\right)}_{\text{advection }v\partial_{x}f^{s}}$ (5)
$+\underbrace{\frac{q^{s}}{m^{s}\alpha^{s}}E(x,t)\left(-\sqrt{\frac{n}{2}}C_{n-1}^{s}(x,t)+\sqrt{\frac{n+1}{2}}C_{n+1}^{s}(x,t)\right)}_{{\text{acceleration }q^{s}/m^{s}E\partial_{v}f^{s}}}=0,$
@@
Note that the SW Hermite closure introduces an error in both the advection $(v\partial_{x}f^{s})$ and acceleration $(q^{s}/m^{s}E\partial_{v}f^{s})$ terms, as highlighted in Eq. (5). The Poisson equation (2) after Galerkin projection is given by
@@
The integral $\mathcal{I}^{0}_{n}\in\mathbb{R}$ recursive properties are as follows
$\mathcal{I}^{0}_{0}=\sqrt{2}\pi^{\frac{1}{4}},\qquad\mathcal{I}_{1}^{0}=0,\qquad\mathrm{and}\qquad\mathcal{I}_{n}^{0}=\sqrt{\frac{n-1}{n}}\mathcal{I}_{n-2}^{0}\qquad\mathrm{for}\qquad n\geq 2.$ (8)
@@ 3 Conservative Closures for the Symmetrically Weighted Hermite Formulation
The main question we seek to answer is: Is there a closure in the form of Eq. (6) that improves the mass, momentum, energy, or Casimir (e.g., $\mathcal{L}_{2}$ norm) conservation properties of the SW Hermite formulation? We begin by deriving the mass, momentum, and energy fluid equations and their respective drift rates with an arbitrary closure, see [19, §4.1] for a more detailed derivation for the closure by truncation. Then, we set the drift rates to zero and derive a conservative closure. The proposed closures are (1) closure by truncation, (2) closure by momentum, and (3) closure by energy. The closures result in a system of partial differential-algebraic equations for the coefficients motivated by the fluid equations, which are the first three moments with respect to the velocity monomials $,1,v,v^{2},$ of the Vlasov equation (1).
@@ 3.1 Mass Conservative Closure (Equivalent to Closure by Truncation)
The integral $\mathcal{I}_{n}^{1}\in\mathbb{R}$ is related to the integral $\mathcal{I}_{n}^{0}\in\mathbb{R}$ in Eq. (8) by
$\mathcal{I}_{n}^{1}=\begin{cases}\alpha^{s}\left(\sqrt{\frac{n+1}{2}}\mathcal{I}^{0}_{n+1}+\sqrt{\frac{n}{2}}\mathcal{I}^{0}_{n-1}\right)=\alpha^{s}\sqrt{2n}\mathcal{I}^{0}_{n-1}=\alpha^{s}\sqrt{2(n+1)}\mathcal{I}_{n+1}^{0}\qquad&\text{if }n\geq 1\text{ and odd,}\\ u^{s}\mathcal{I}^{0}_{n}\qquad&\text{if }n\geq 0\text{ and even.}\end{cases}$
@@
The closure cannot impact the mass drift rate if $N_{v}$ is even. However, if $N_{v}$ is odd, then setting $\int_{0}^{\ell}E(x,t)C_{N_{v}}^{s}(x,t)\mathrm{d}x=0$ causes the total mass drift rate to vanish. The closure by truncation, i.e. $C_{N_{v}}^{s}(x,t)=0$, is one example of a closure resulting in a total mass conservative formulation and the only closure that satisfies the continuity equation.
@@ 3.2 Momentum Conservative Closure
The integral $\mathcal{I}_{n}^{2}\in\mathbb{R}$ is related to the integral $\mathcal{I}_{n}^{0}\in\mathbb{R}$ in Eq. (8) by
$\mathcal{I}_{n}^{2}=\begin{cases}2\alpha^{s}u^{s}\left(\sqrt{\frac{n+1}{2}}\mathcal{I}^{0}_{n+1}+\sqrt{\frac{n}{2}}\mathcal{I}^{0}_{n-1}\right)=2u^{s}\mathcal{I}_{n}^{1}\qquad&\text{if }n\geq 1\text{ and odd,}\\ (\alpha^{s})^{2}\left(\sqrt{\frac{(n+1)(n+2)}{4}}\mathcal{I}^{0}_{n+2}+\left(\frac{2n+1}{2}+\left(\frac{u^{s}}{\alpha^{s}}\right)^{2}\right)\mathcal{I}^{0}_{n}+\sqrt{\frac{n(n-1)}{4}}\mathcal{I}^{0}_{n-2}\right)\qquad&\text{if }n\geq 0\text{ and even.}\end{cases}$
@@
Therefore, the fluid momentum equation is satisfied and the momentum drift rate vanishes if
$C_{N_{v}}^{s}(x,t)=-C_{N_{v}-1}^{s}(x,t)\begin{cases}\frac{\alpha^{s}\sqrt{2N_{v}}}{u^{s}}\qquad&\text{if }N_{v}\text{ is odd},\\ \frac{u^{s}}{\alpha^{s}\sqrt{2N_{v}}}\qquad&\text{if }N_{v}\text{ is even}.\\ \end{cases}$ (16)
We refer to this closure as closure by momentum. For both cases of $N_{v}$ being odd or even, the closure term $C_{N_{v}}^{s}(x,t)$ is proportional to the last Hermite expansion coefficient $C_{N_{v}-1}^{s}(x,t)$. It is also important to mention that when $N_{v}$ is odd this closure requires the velocity shifting parameter $u^{s}\neq 0$ for all species $s$, which can be easily imposed by performing a Galilean transformation. In the case of $N_{v}$ being even and $u^{s}=0$, the closure by truncation also satisfies the momentum equation and thus preserves the total momentum.
@@ 3.3 Energy Conservative Closure
where $\mu^{s}_{n}(x,t)=\frac{q^{s}}{m^{s}}\int_{0}^{x}E(y,t)C_{n}^{s}(y,t)\mathrm{d}y$. The extra term involving $\mu_{n}^{s}(x,t)$ stems from integrating the SW Poisson equation (7) to arrive at the SW Ampere’s equation. The total energy drift rate in Eq. (18) is zero, if
@@ 4 Closure Structure Preservation: Hyperbolicity and Anti-symmetry
where $\sigma^{s}\in\mathbb{R}$ linearly relates the closure term $C_{N_{v}}^{s}(x,t)$ to $C_{N_{v}-1}^{s}(x,t)$. Similarly, by Eq. (5), the source operator with a general closure in the form of Eq. (6) is given by
@@ 4.1 Hyperbolicity
So far we have treated the case with constant $\alpha^{s}$ and $u^{s}$. However, there is a strong mathematical and physical motivation to adapt spatially and temporally the Hermite shifting and scaling parameters, denoted as $u^{s}$ and $\alpha^{s}$ in Eq. (4). The mathematical considerations stem from improving the convergence of the Hermite spectral discretization. The physical motivation stems from the Hermite basis function in Eq. (4) resembling a Maxwellian distribution (thermodynamic equilibrium), such that $u^{s}$ mimics the characteristic mean flow and $\alpha^{s}$ mimics the thermal velocity. Therefore, the Hermite parameters can be adapted in space and time based on the evolution of such macroscopic quantities. An adaptive approach based on evolving the macroscopic quantities was first proposed in the seminal paper by Grad [13], which suffers from a lack of hyperbolicity. On the other hand, recent work by [29] showed that the adaptivity in time of the AW Hermite parameters (i.e. $u^{s}$ and $\alpha^{s}$) does not alter the governing equations of the expansion coefficients and relies on projecting the expansion coefficients at a fixed instant in time of the Hermite parameters. This allows the adaptive in-time AW and SW formulations to remain hyperbolic. The adaptivity in space of the Hermite parameters introduces new terms in the form of $C_{n}^{s}(x,t)\frac{\partial}{\partial x}\alpha^{s}(x)$ and $C_{n}^{s}(x,t)\frac{\partial}{\partial x}u^{s}(x)$ as additional “source” terms. These terms do not directly impacts the Jacobian flux matrix and similar to time-adaptivity require a Hermite projection step onto the new basis at a fixed instant in time [24]. Therefore, we anticipate that by following the proposed adaptive approach for the AW Hermite method in [29, 24], a space and time adaptive approach of the SW Hermite method also preserves the system’s hyperbolicity.
@@ 4.2 Anti-symmetry
Anti-symmetry (also known as skew-symmetry) is an additional important structure for a spectral method to preserve. Aside from capturing the correct dynamics, an anti-symmetric discretization guarantees unconditional numerical stability [8, 19]. Our previous work [19, §3.2] explains in detail the anti-symmetric structure of the Vlasov advection operator. In short, it requires the flux Jacobian matrix in Eq. (21) to be symmetric, i.e. $\mathbf{J}^{s}=(\mathbf{J}^{s})^{\top}$, and the source operator in Eq. (22) to be anti-symmetric, i.e. $\mathbf{S}^{s}(x,t)=-\mathbf{S}^{s}(x,t)^{\top}$. As previously stated in [19, §3.2.2], anti-symmetry is only preserved via the closure by truncation, otherwise $\mathbf{S}^{s}(x,t)\neq-\mathbf{S}^{s}(x,t)^{\top}$. The violation of anti-symmetry stems from $\beta^{s}(x,t)\neq 0$ in Eq. (22). Additionally, the adaptivity in time of the Hermite parameters following the work by [29] does not change the governing equations and will preserve anti-symmetry for the closure by truncation. However, adaptivity in space breaks anti-symmetry by introducing non-zero terms on the main diagonal of the source operator.
@@ 5 Numerical Results
How does the choice of closure influence the conservation properties? Figure 1 shows the total mass, momentum, energy, and $\mathcal{L}_{2}$ norm conservation results with $N_{v}=\{10,11\}$ for the three proposed closures: (1) closure by truncation, (2) closure by momentum, and (3) closure by energy. The numerical results agree with the analytical derivations in section 3. Closure by truncation can conserve the $\mathcal{L}_{2}$ norm of the particle distribution function and total mass if $N_{v}$ is odd, closure by momentum can only conserve total momentum, and closure by energy can only conserve total energy. Figure 1 indicates that the magnitude of the drift rate is comparable for the three closures. Additionally, the conservation relative error fluctuates around zero for mass, momentum, and energy, and is mainly positive for the $\mathcal{L}_{2}$ norm. The closure by momentum and energy with $N_{v}=10$ becomes unstable before the final timestamp. Recall that numerical stability properties are guaranteed only for the closure by truncation since it preserves anti-symmetry.
How does the choice of closure impact the electric field? Figure 2 compares the electric field with $N_{v}=11$ Hermite discretization for the three closures. We compare the simulation results with a high-resolution simulation using $N_{v}=1001$ and closure by truncation. To ensure that the choice of closure does not influence the high-resolution simulation, we verify that the magnitude of the last Hermite coefficient stays close to machine precision throughout the entire simulation. The results for the closure by momentum and closure by energy with $N_{v}=11$ are polluted by high wavenumber oscillations in comparison to the closure by truncation. This suggests that the closure by truncation may be the most appropriate conservative closure for the SW Hermite method.
(a) $\mathcal{L}_{2}$ conservation
(b) mass conservation
(c) momentum conservation
(d) energy conservation
Figure 1: Conservation relative error of total (a) $\mathcal{L}_{2}$ norm, (b) mass, (c) momentum, and (d) energy for the Langmuir wave test. The numerical results verify the analytic derivations in Eqns. (12), (15), and (18).